OpenStudy (sasogeek):

The annual rent of rectangular plot of land is $1,024,000 at a rate of $320 per square meter. If the length is 80m, find the width of plot.

OpenStudy (anonymous):

give m e 1 sec

OpenStudy (ldmoriarty):

K, You must first find the total area of the plot of land. You are given the total rent and the rent per square meter. Divide the total rent by the rent per sq meter to find the total area in square meters. Then divide the area by the length to find the width. I get 40 meters. Anyone? Am I right or wrong here?

OpenStudy (anonymous):

$1,024,000=320x x = 1,024,000/320 x =3.200 80*w=3.200 w= 3.200/80 w=40

OpenStudy (anonymous):

A=3.200

OpenStudy (anonymous):

Medal help?